# Gradient-Free Adaptive Global Pruning for Pre-trained Language Models

### v1

@@ 5 Experiments
Sparsity
Method
BoolQ
RTE
HellaSwag
WinoGrande
ARC-e
ARC-c
OBQA
Mean
Dense
70.46
61.82
54.27
69.02
70.47
35.49
30.20
55.96
70%
SparseGPT
68.78
58.48
53.83
67.64
69.15
34.30
29.60
54.54
AdaGP
69.11
61.73
53.97
68.43
69.78
34.73
29.80
55.36
80%
SparseGPT
64.86
60.65
49.73
61.40
61.91
31.74
24.20
50.64
AdaGP
65.41
59.57
50.65
61.96
62.71
32.25
26.50
51.29
90%
SparseGPT
37.83
53.79
25.96
49.88
26.47
20.22
12.60
32.39
AdaGP
43.55
52.35
26.32
50.04
27.31
20.56
14.00
33.45
2:4
SparseGPT
69.54
58.12
54.08
67.25
69.95
34.81
29.80
54.79
AdaGP
69.67
58.12
54.08
68.88
70.45
35.49
30.20
55.27
Comparison Methods.
We compare against the standard magnitude pruning baseline Zhu and Gupta (2017) applied locally. We also compare with two state-of-the-art local pruning methods, SparseGPT Frantar and Alistarh (2023) and Wanda Sun et al. (2023).
@@ 5.1 Results and Analyses
We begin by exploring the pruning capabilities of AdaGP across various model sizes in comparison to the baseline methods. For each model, we consider unstructured sparsity ranging from 70% to 90% with a 10% increment, as well as a full 2:4 semi-structured sparsity. We conduct a sensitivity study on the calibration sample sizes (see Appendix A.4) and use calibration sample sizes between 32 and 64 for all experiments. Moreover, except for the OPT-125m model, we prune the first 50% of the transformer decoder layers in each model to achieve a balance between the computation resources and showcasing the performances. Detailed results can be found in Table 1 and Table 3 (see Appendix A.1). The perplexity results on the three datasets for the un-pruned dense models are reported next to the names of the models. Note that we exclude the results for OPT-350m and adjust 90% unstructured sparsity for OPT-1.3b to 85% unstructured sparsity for numerical stability. We provide a discussion on the selection of hyperparameters of $\alpha$ and $\beta$ (used in Eq. 5) in Appendix A.5.
@@
From the table, it shows that perplexity increases with increasing sparsity. Moreover, we observe a trend of decreasing perplexity for SparseGPT and AdaGP at the same sparsity with increasing model sizes. However, such a trend is not obvious for Magnitude and Wanda. We also observe that SparseGPT and AdaGP consistently outperform Magnitude and Wanda by a significant margin. For smaller sparsity, AdaGP achieves comparable perplexity to SparseGPT. As we increase the sparsity, AdaGP starts to demonstrate noticeable improvements over SparseGPT. In numerous instances, AdaGP achieves perplexity reductions of more than 50% compared to SparseGPT.
Computation Time vs. Model Sizes.
Next, we study how the computation time for a single epoch of SparseGPT and AdaGP varies with different model sizes, as illustrated in Figure 3. We observe that the time taken for a single epoch increases with increasing model sizes. The rate at which the time taken increases is comparable for SparseGPT and AdaGP. Additionally, although SparseGPT executes only one epoch for each layer to prune, whereas AdaGP may execute multiple epochs depending on the parameter configuration, we find that in most cases, few epochs (less than 10) of AdaGP suffice. This suggests that the proposed AdaGP remains computationally efficient.
Figure 3: Time for one step of SparseGPT and AdaGP.
Zero-shot Experiments.
@@
To conclude the evaluations and discussions, we show the results for several zero-shot tasks in Table 2 comparing SparseGPT and AdaGP and three model sizes. These evaluations are known to be relatively noisy Dettmers et al. (2022), but more interpretable. We also report the results for zero-shot tasks from the un-pruned dense models in the “Dense" row. We see that the accuracy of both methods decreases with increasing sparsity, which is expected, as more parameters are pruned. A similar trend of increasing accuracy with increasing model size is observed too. Across all the tasks, OBQA and ARC-c remain the most challenging ones as the accuracy for both methods remain around or below 30% while both methods perform well for BoolQ, RTE, WinoGrande, and ARC-e. In general, AdaGP is able to achieve higher accuracy in almost all tasks across the three models of different sizes.
@@ 6 Conclusion
Our work presents AdaGP, a cutting-edge framework poised to redefine the compression of LLMs through sparsity. By adeptly circumventing the scalability issues of global pruning and addressing the local suboptimality of existing methods, AdaGP stands as a significant advancement in the field. Our empirical results affirm its efficacy, particularly in high-sparsity environments, where it achieves a notable reduction in perplexity, thereby setting a new precedent for model compression. The versatility and minimal computational overhead of AdaGP complement its integration with current pruning technologies, underscoring its potential as a universal tool for enhancing the performance and accessibility of LLMs.
7 Limitations
While AdaGP marks a significant step forward in efficient pruning of large language models, it is important to acknowledge the inherent trade-off associated with any model compression technique. Firstly, while our method reduces the complexity of LLMs and enhances computational efficiency, there is an inevitable balance between sparsity and performance that requires careful calibration. Additionally, the effectiveness of AdaGP, similar to any pruning method, may vary across different models and tasks, and its generalizability to all scenarios remains an area for further exploration. Lastly, our approach assumes certain structural properties of the neural network, such as layer-wise decomposability, which may not hold for all architectures. These limitations, while not undermining the contributions of AdaGP, suggest avenues for future research and refinement.
Acknowledgement
This research used resources of the Argonne Leadership Computing Facility at Argonne National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under contract DE-AC02-06CH11357. We gratefully acknowledge the computing resources provided on Swing, a high-performance computing cluster operated by the Laboratory Computing Resource Center at Argonne National Laboratory.
Government license
The submitted manuscript has been created by UChicago Argonne, LLC, Operator of Argonne National Laboratory ("Argonne”). Argonne, a U.S. Department of Energy Office of Science laboratory, is operated under Contract No. DE-AC02-06CH11357. The U.S. Government retains for itself, and others acting on its behalf, a paid-up nonexclusive, irrevocable worldwide license in said article to reproduce, prepare derivative works, distribute copies to the public, and perform publicly and display publicly, by or on behalf of the Government. The Department of Energy will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (http://energy.gov/downloads/doe-public-access-plan)
@@ Appendix A Appendix
A.1 OPT-125m
Detailed results for OPT-125m are in Table 3. For a thorough investigation, we additionally consider unstructured sparsity of 50% and 60% as well as pruning the full model (instead of the first 50% of the transformer decoder layers). Similar to other models, we report the perplexity results for the un-pruned dense model next to the name of the model in the table. In particular, we see that SparseGPT and AdaGP outperform Magnitude and Wanda with a significant margin across different level of sparsity. AdaGP achieves similar perplexity with SparseGPT at smaller sparsity, but demonstrates much better perplexity for larger sparsity. Similar perplexity trends are observed across the three datasets, although, PTB, having the highest perplexity for each sparsity and method, seems to be the most challenging dataset among the three.
OPT-125m (WikiText2: 27.66; PTB: 38.99; C4: 26.56)
Sparsity Method WikiText2 PTB C4 Sparsity Method WikiText2 PTB C4
50% Magnitude 193.39 276.33 141.03 80% Magnitude 4890.96 4121.49 3213.85
Wanda 39.58 57.44 35.12 Wanda 1912.45 2512.93 1066.86
SparseGPT 37.49 55.91 33.63 SparseGPT 2072.12 1952.85 1050.83
AdaGP 37.09 55.82 33.40 AdaGP 1358.10 1418.09 654.54
60% Magnitude 920.29 1146.49 598.45 90% Magnitude 6613.18 5380.80 4475.29
Wanda 76.30 106.92 61.33 Wanda 4940.89 4337.27 3126.02
SparseGPT 62.22 87.70 49.85 SparseGPT 6131.57 6963.27 2443.33
AdaGP 59.28 84.36 47.78 AdaGP 5291.64 5067.41 2003.09
70% Magnitude 3806.96 3429.35 2263.37 2:4 Magnitude 342.51 798.99 223.19
Wanda 351.83 412.52 248.94 Wanda 83.81 113.01 65.78
SparseGPT 239.26 265.83 156.33 SparseGPT 61.44 92.30 51.42
AdaGP 208.46 255.75 137.72 AdaGP 56.55 81.46 46.14
Table 3: Perplexity on pruned OPT-125m; the lower the perplexity, the better
A.2 Pseudo-code of AdaGP
Input: Pre-trained weight matrices $\mathbf{W}_{\ell}$ and $\mathbf{W}_{\ell+1}$, inputs $\boldsymbol{a}_{\ell-1}^{pre}$ and $\boldsymbol{a}_{\ell}^{pre}$, and outputs $\mathbf{z}_{\ell}^{pre}$ and $\mathbf{z}_{\ell+1}^{pre}$ for FFN’s up-scaling and down-scaling layers $\ell$ and $\ell+1$; target sparsity $\rho$; constraint weight hyperparameters $\alpha$, $\beta$.
1AdaGP():
@@ A.2 Pseudo-code of AdaGP
5             $\mathbf{W}_{\ell}=\mathbf{z}_{\ell}\boldsymbol{a}_{\ell-1}^{\dagger}$, $\mathbf{W}_{\ell+1}=\mathbf{z}_{\ell+1}^{pre}\boldsymbol{a}_{\ell}^{\dagger}$
@@
8             $\mathbf{W}_{\ell+1}=\mathbf{M}_{\ell+1}\odot\widehat{\mathbf{W}}_{\ell+1}$, $\mathbf{W}_{\ell}=\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell}$
@@ 16                        ${(\mathbf{z}_{\ell})}_{j}={(\mathbf{z}_{\ell}^{(2)})}_{j}$triangleright Update outputs
Algorithm 1 Adaptive Global Pruning for the FFN module of an LLM.
@@
The Adaptive Global Pruning (AdaGP) algorithm presented in Algorithm 1 focuses on the optimization of the Feed-Forward Network (FFN) module within a Large Language Model (LLM). The key inputs to the algorithm include the pre-trained weight matrices, inputs and outputs for both the up-scaling and down-scaling linear layers of the FFN, along with a set of hyperparameters and a target sparsity. The goal of AdaGP is to achieve a targeted level of sparsity in the linear layers without significantly compromising the model’s performance.
@@
Initiating with the pre-trained weights, AdaGP employs a series of pruning and activation update steps across $K$ epochs. In each epoch, it solves optimization problems to prune the weights in both up-scaling and down-scaling layers, followed by updating the activation variables. The utilization of SparseGPT solver for pruning and the strategic update of activations ensure that the pruned network behaves as similarly as possible to that of the original network. The final output of the algorithm is a pair of pruned weight matrices for the consecutive layers, which are expected to deliver comparable, if not improved, performance with a reduced number of non-zero parameters.
Figure 4: Visualization of the AdaGP Algorithm Applied to the FFN Module in a transformer decoder. This diagram illustrates the flow of data through the layers of the decoder, where AdaGP performs global pruning to enhance sparsity and computational efficiency without compromising the model’s performances. Each decoder layer is shown with its core components: Masked Multi-Head Attention and Feed Forward network, with AdaGP focusing on the latter for pruning. For other linear layers outside of the FFN modules, AdaGP follows the same local pruning as SparseGPT does to prune them.
A.3 Visualization of AdaGP for FFN Global Pruning
Figure 4 presents a schematic representation of the AdaGP algorithm applied to the Feed-Forward Network (FFN) module within a transformer model’s decoder layer. This visualization highlights the process of global pruning applied to the FFN module, aimed at reducing the model’s complexity while preserving its ability to capture essential patterns in the data.
The AdaGP method operates within the decoder block, selectively pruning the weights of the FFN module. Each decoder block comprises a series of operations. It starts with masked multi-head attention and follows by additive normalization. Subsequently, it passes through the feed-forward network in which the pruning happens with AdaGP. The figure illustrates the flow of information from an input token through successive layers of the decoder. AdaGP not only enhances computational efficiency but also has the potential to improve model generalization by eliminating redundant parameters.
A.4 Calibration Samples
@@
Figure 5 presents how perplexity changes with the calibration sample sizes on the datasets PTB and C4, respectively. On one hand, in both plots, as the number of calibration samples increases, the perplexity decreases for both SparseGPT and AdaGP. This suggests that having more calibration samples can be beneficial in the pruning process. Perplexity decreases more rapidly from 8 samples to 32 samples. Beyond 32 samples, the rate at which perplexity decreases starts to slow down. For PTB dataset, the perplexity even increases from 64 samples to 128 samples. On the other hand, increasing the number of calibration samples requires more computational resources, e.g., memory and computation time, in the execution. A calibration sample size should be between 32 and 64 can therefore strike a balance between good performances and computational efficiency. Lastly, we want to point out that the figures show that AdaGP starts to achieve better perplexity than SparseGPT does from 32 samples and remains better for 64 and 128 samples.
@@
A.5 Hyperparameter $\alpha$ and $\beta$ Selection
Hyperparameters $\alpha$ and $\beta$ are used in Eq. 5. We select $\alpha$ and $\beta$ from the set $\{0.1,1.0,5.0\}$ and perform a study on model OPT-2.7b to understand the impact of the hyperparameters. We determine that $\alpha=\beta=0.1$ give the best and numerically stable performances.

### v3

@@ 5 Experiments
Sparsity Method BoolQ RTE HellaSwag WinoGrande ARC-e ARC-c OBQA Mean
Dense 70.46 61.82 54.27 69.02 70.47 35.49 30.20 55.96
70% SparseGPT 68.78 58.48 53.83 67.64 69.15 34.30 29.60 54.54
SparseLLM 69.11 61.73 53.97 68.43 69.78 34.73 29.80 55.36
80% SparseGPT 64.86 60.65 49.73 61.40 61.91 31.74 24.20 50.64
SparseLLM 65.41 59.57 50.65 61.96 62.71 32.25 26.50 51.29
90% SparseGPT 37.83 53.79 25.96 49.88 26.47 20.22 12.60 32.39
SparseLLM 43.55 52.35 26.32 50.04 27.31 20.56 14.00 33.45
3:4 SparseGPT 55.81 51.26 33.64 54.54 42.05 21.33 21.00 39.95
SparseLLM 60.83 54.15 39.35 55.41 45.24 24.06 22.20 43.03
LlaMA-2 7b
Sparsity Method BoolQ RTE HellaSwag WinoGrande ARC-e ARC-c OBQA Mean
Dense 75.05 66.43 56.92 69.93 75.34 41.89 34.40 59.99
70% SparseGPT 68.26 57.04 39.67 59.04 60.9 28.58 20.60 47.73
SparseLLM 67.61 57.31 40.12 61.39 59.39 28.76 21.40 48.13
80% SparseGPT 59.36 52.71 28.83 48.7 34.22 18.34 14.40 36.65
SparseLLM 60.12 53.07 28.62 50.59 34.55 18.69 14.30 37.13
90% SparseGPT 39.02 52.34 26.66 47.80 28.32 17.37 12.40 31.99
SparseLLM 39.45 52.71 26.79 51.17 28.32 19.52 12.50 32.92
3:4 SparseGPT 53.94 54.15 28.09 49.17 31.57 17.41 14.80 35.59
SparseLLM 57.34 53.43 28.26 48.86 32.45 18.17 14.4 36.13
LlaMA-2 13b
Sparsity Method BoolQ RTE HellaSwag WinoGrande ARC-e ARC-c OBQA Mean
Dense 77.89 70.40 59.94 72.77 77.40 46.50 33.20 62.59
70% SparseGPT 70.03 53.43 42.20 66.54 64.94 31.66 25.40 50.60
SparseLLM 69.87 54.15 42.50 68.64 64.97 31.40 25.80 51.05
80% SparseGPT 62.69 52.71 28.94 50.91 36.24 18.17 14.00 37.67
SparseLLM 64.39 52.86 29.19 51.46 35.69 18.77 14.20 38.08
90% SparseGPT 50.21 51.35 26.71 49.14 26.68 19.71 13.2 33.86
SparseLLM 55.35 52.05 26.89 51.34 27.35 19.62 14.20 35.26
3:4 SparseGPT 61.28 53.71 28.40 47.99 33.21 18.26 14.00 36.69
SparseLLM 61.71 55.71 28.56 51.62 32.11 18.49 13.8 37.43
Comparison methods. We compare against three baselines, magnitude pruning Zhu and Gupta (2017) applied locally, and two other state-of-the-art local pruning methods, SparseGPT Frantar and Alistarh (2023) and Wanda Sun et al. (2023).
@@ 5.1 Results and analyses
Pruning vs. model sizes. We begin by exploring the pruning capabilities of SparseLLM across various model sizes in comparison to baseline methods. For each model, we consider unstructured sparsity ranging from 70% to 90% with a 10% increment, as well as a 3:4 semi-structured sparsity. The 3:4 semi-structured sparsity is inspired by our preliminary results that suggest good performance SparseLLM at high sparsity regimes. However, note that two of our baselines, Magnitude and Wanda, are unable to be configured to this sparsity out-of-box. We conduct a sensitivity study on the calibration sample sizes (see Appendix A.3) and use calibration sample sizes between 32 and 64 for all experiments. Moreover, we prune the first 50% of the Transformer decoder layers in each model to achieve a balance between the computation resources and the performances. Detailed results can be found in Table 1 and Table 5 (see Appendix A.5). The perplexity results of the dense models are reported next to the names of the models.
Figure 3: Fast convergence of SparseLLM. Training loss per epoch for pruning layer 3 of OPT-125m at 80% sparsity (Left) and layer 6 of LlaMA-2 13b at 70% sparsity (Right).
@@
From the table, it shows a general trend of increasing perplexity with increasing sparsity. Moreover, we observe a trend of decreasing perplexity for SparseGPT and SparseLLM at the same sparsity with increasing model sizes. However, such a trend is not obvious for Magnitude and Wanda. We also observe that SparseGPT and SparseLLM consistently outperform Magnitude and Wanda by a significant margin. For smaller sparsity, SparseLLM achieves comparable perplexity to SparseGPT. As we increase the sparsity, SparseLLM starts to demonstrate noticeable improvements over SparseGPT. In numerous instances for the OPT model family, SparseLLM achieves perplexity reductions of more than 50% compared to SparseGPT. We also see that performance improvements from SparseLLM over SparseGPT are more significant for the OPT model family than the LlaMA-2 model family.
@@
Zero-shot experiments. To further conclude the evaluations and discussions, we show results for several zero-shot tasks in Table 2 and Table 6 (see Appendix A.5) comparing SparseGPT and SparseLLM. These evaluations are known to be relatively noisy Dettmers et al. (2022), but more interpretable. We also report the results for zero-shot tasks from the dense models in the “Dense" row. We see that the accuracy of both methods decreases with increasing sparsity, which is expected, as more parameters are pruned. A similar trend of increasing accuracy with increasing model size is observed too. Across all the tasks, OBQA and ARC-c remain the most challenging ones as the accuracy for both methods is 30% or below 30% while both methods perform well for BoolQ, RTE, WinoGrande, and ARC-e. In general, SparseLLM is able to achieve higher accuracy in the majority of tasks across the models of different sizes in both OPT and LlaMA-2 model families.
Training loss vs. epochs in SparseLLM. Figure 3 illustrates the change in training loss over epochs for SparseLLM, with the training loss plotted on a scale of $10^{3}$ for clarity. We observe that the training loss decreases rapidly during the initial epochs, highlighting the efficiency of SparseLLM in achieving effective global pruning within a short period. This rapid convergence is largely due to the closed-form solutions employed by SparseLLM for various subproblems, which streamline the pruning process and ensure optimal layer-wise pruning without extensive iterative computations. These analytical solutions enable SparseLLM to perform precise pruning operations quickly, making it a powerful tool for optimizing large-scale models like LlaMA, significantly reducing model size while maintaining high accuracy.
@@ 6 Conclusion
Our work presents SparseLLM, a cutting-edge framework poised to redefine the compression of LLMs through sparsity. By adeptly circumventing the scalability issues of global pruning and optimizing the local suboptimality of existing methods, SparseLLM stands as a significant advancement in the field. Our empirical results affirm its efficacy, particularly in high-sparsity environments, where it achieves a notable reduction in perplexity, thereby setting a new precedent for model compression. The versatility and minimal computational overhead of SparseLLM complement its integration with current pruning technologies, underscoring its potential as a universal tool for enhancing the performance and accessibility of LLMs.
@@ Appendix A Appendix
This section includes supplemental materials (pseudo-code, additional experiments, and plots).
A.1 Pseudo-code of SparseLLM
1
Input: An OPT decoder layer containing FFN and MHA modules. FFN’s up-scaling linear layer pre-trained weight matrix $\mathbf{W}_{\ell}$, FFN’s down-scaling linear layer pre-trained weight matrix $\mathbf{W}_{\ell+1}$, input of the up-scaling linear layer $\boldsymbol{a}_{\ell-1}^{pre}$, output of the down-scaling linear layer $\mathbf{z}_{\ell+1}^{pre}$, target sparsity $\rho$, constraint weight hyperparameters $\alpha$, $\beta$.
2 SparseLLM on FFN():
3
@@ 2 SparseLLM on FFN():
7
8             $\mathbf{W}_{\ell}=\mathbf{z}_{\ell}\boldsymbol{a}_{\ell-1}^{\dagger}$,
9             $\mathbf{W}_{\ell+1}=\mathbf{z}_{\ell+1}\boldsymbol{a}_{\ell}^{\dagger}$
@@
12             $\mathbf{W}_{\ell+1}=\mathbf{M}_{\ell+1}\odot\widehat{\mathbf{W}}_{\ell+1}$,
13             $\mathbf{W}_{\ell}=\mathbf{M}_{\ell}\odot\widehat{\mathbf{W}}_{\ell}$
@@ 21                        ${(\mathbf{z}_{\ell})}_{j}={(\mathbf{z}_{\ell}^{(2)})}_{j}$triangleright Update outputs
26
27 SparseLLM on MHA():
28       for each linear layer $\ell$ in MHA module do
29
30             Fix $\mathbf{z}_{\ell}=\mathbf{z}_{\ell}^{pre}$, $\boldsymbol{a}_{\ell}=\boldsymbol{a}_{\ell}^{pre}$triangleright Fix intermediate variables
32
33
34      return $\mathbf{W}_{\ell}$} for all linear layers in MHA
35
Algorithm 1 SparseLLM Pruning of OPT Models.
@@ 27 SparseLLM on MHA():
The SparseLLM algorithm presented in Algorithm 1 demonstrates how SparseLLM works on an OPT decoder layer. The key inputs to the algorithm include the pre-trained weight matrices for both the up-scaling and down-scaling linear layers of the FFN, along with a set of hyperparameters and a sparsity constraint. The goal of SparseLLM is to achieve a targeted level of sparsity in the linear layers without significantly compromising the model’s performance.
@@
Initiating with the pre-trained weights, SparseLLM employs a series of pruning and activation update steps across $K$ iterations. In each iteration, it solves optimization problems to prune the current and subsequent layer weights, followed by updating the activation variables. The utilization of SparseGPT solvers for pruning and the strategic update of activations ensures that the pruned network approximates the original network’s behavior as closely as possible. The final output of the algorithm is a pair of pruned weight matrices for the consecutive layers, which are expected to deliver comparable or improved performance with a reduced number of parameters.
A.2 Two-layer Demo on the Details behind our Global Pruning
Figure 4 illustrates the SparseLLM pruning method compared to conventional global pruning and local pruning, using a two-layer neural network as an abstraction for simplicity. The figure is divided into three main parts:
On the left, conventional global pruning is depicted. This method applies a global mask to the entire network, resulting in significant memory costs due to poor scalability. Both functions $f_{1}$ and $f_{2}$ are pruned using the same mask across all layers, leading to high memory usage.
In the middle, local pruning is shown, where each layer is pruned independently. This approach reduces memory costs by applying separate masks to each layer. However, it inevitably sacrifices performance because it ignores global supervision, which can lead to suboptimal pruning decisions that do not consider the network as a whole.
On the right, the adaptive global pruning method of SparseLLM is presented. This method achieves global pruning with low memory cost by leveraging auxiliary variables and soft constraints. It combines the benefits of global pruning—considering the entire network structure—with efficient memory usage. The introduction of auxiliary variables allows for flexible and adaptive pruning, ensuring that the overall performance of the network is maintained while keeping memory costs low.
Thus, the figure highlights the trade-offs between different pruning strategies. Conventional global pruning incurs high memory costs, local pruning reduces memory usage at the expense of performance, and the adaptive global pruning of SparseLLM strikes a balance by maintaining performance with lower memory requirements through the use of auxiliary variables and soft constraints.
Figure 4: Illustration of SparseLLM pruning method compared to conventional global pruning and local pruning. We consider a two-layer neural network as an abstraction for simplicity. Global pruning (left) is memory prohibitive due to poor scalability. Local pruning (mid) considers pruning each layer independently, while inevitably sacrificing performance due to the ignorance of global supervision. Our adaptive global pruning (right) achieves global pruning with low memory cost by leveraging auxiliary variables and soft constraints.
A.3 Calibration Samples
@@
Figure 5 and Figure 6 present how perplexity changes with the calibration sample sizes on the datasets PTB and C4 for OPT-2.7b and LlaMA-2 7B, respectively. In both figures, as the number of calibration samples increases, the perplexity decreases for both SparseGPT and SparseLLM. This indicates that having more calibration samples can be beneficial in the pruning process. Perplexity decreases more rapidly from 8 samples to 32 samples. Beyond 32 samples, the rate at which perplexity decreases starts to slow down. In addition, increasing the number of calibration samples requires more computational resources, e.g., memory and computation time, in the overall pruning process. This suggests that the calibration sample sizes should be between 32 and 64 to ensure good performance while maintaining computational efficiency. Lastly, the figures show that SparseLLM achieves better perplexity than SparseGPT does with 32 or larger sample sizes for both OPT and LlaMA-2 models.
@@
Figure 6: Sensitivity of LlaMA-2 7b models on the calibration sample sizes for datasets PTB and C4.
A.4 Computation Time vs. Model Sizes
We study how the computation time of SparseGPT and SparseLLM varies with different model sizes, as illustrated in Table 3 and Table 4 for OPT models and LlaMA-2 models. The rate at which the time taken increases is comparable for SparseGPT and SparseLLM as the model size increases. Additionally, computation time for SparseLLM are reported for a configuration of 4 to 10 epochs. As we have reported in Section 5, SparseLLM can reduce the training loss in as few as 2 to 3 epochs. This suggests that the proposed SparseLLM remains computationally efficient.
Table 3: Computation time in seconds of OPT models.
Method OPT-125m OPT-1.3b OPT-2.7b OPT-6.7b OPT-13b OPT-30b OPT-66b
SparseGPT 1.47 4.03 4.36 8.15 10.46 20.94 34.78
SparseLLM 19.58 74.76 97.44 184.96 209.26 164.34 270.38
Table 4: Computation time in seconds of LlaMA-2 models.
Method Llama-2 7b Llama-2 13b
SparseGPT 179.08 314.95
SparseLLM 2201.99 4797.07
A.5 Experiment Results for Additional Models
Detailed results on perplexity and zero-shot task accuracy for additional models are reported in Table 5 and Table 6. Similar to other models, we report the perplexity results for the dense model next to the name of the model in the table. In particular, we see that SparseGPT and SparseLLM outperform Magnitude and Wanda with a significant margin across different sparsity. SparseLLM shares similar perplexity with SparseGPT for smaller sparsity but demonstrates much better perplexity for larger sparsity. Similar perplexity trends are observed across all three datasets, although, PTB, having the highest perplexity for each sparsity and method, is likely the most challenging dataset among the three. For the zero-shot taks accuracy, we see that SparseLLM achieves comparable results to SparseGPT for smaller perplexity and the performance improvements are more obvious and significant with higher sparsity.
Table 5: Perplexity in high sparsity regimes ($\geq 70\%$); the lower the perplexity, the better.
OPT-125m (WikiText2 (WT2): 27.66; PTB: 38.99; C4: 26.56)
Sparsity 70% 80% 90% 3:4
Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
Magnitude 3806.96 3429.35 2263.37 4890.96 4121.49 3213.85 6613.18 5380.80 4475.29 - - -
Wanda 351.83 412.52 248.94 1912.45 2512.93 1066.86 4940.89 4337.27 3126.02 - - -
SparseGPT 239.26 265.83 156.33 2072.12 1952.85 1050.83 6131.57 6963.27 2443.33 1482.61 2215.44 657.26
SparseLLM 208.46 255.75 137.72 1358.10 1418.09 654.54 5291.64 5067.41 2003.09 914.87 1210.84 450.01
OPT-6.7b (WikiText2 (WT2): 10.86; PTB: 15.77; C4: 12.71)
Sparsity 70% 80% 90% 3:4
Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
Magnitude 7054.21 5437.44 4850.25 7937.49 5971.86 6031.54 2.4e4 2.5e4 2.1e4 - - -
Wanda 54.95 129.73 116.67 1493.58 1196.93 996.00 2.1e4 2.0e4 1.8e4 - - -
SparseGPT 12.27 18.90 15.28 31.04 51.26 29.42 8871.24 5713.57 3797.20 570.08 361.81 328.18
SparseLLM 12.16 18.39 14.93 23.96 39.32 26.97 2095.85 1842.48 953.44 83.36 128.99 62.11
OPT-13b (WikiText2 (WT2): 10.13; PTB: 14.52; C4: 12.06)
Sparsity 70% 80% 85% 3:4
Dataset WT2 PTB C4 WT2 PTB C4 WT2 PTB C4 WT2 PTB C4
Magnitude 9037.12 7734.58 5909.47 1.1e4 9140.88 6340.22 1.3e4 1.3e4 9087.50 - - -
Wanda 30.94 39.26 33.31 4216.04 2894.77 2450.57 1.1e4 1.1e4 7244.96 - - -
SparseGPT 10.89 16.35 13.39 21.42 33.62 21.01 8408.03 6380.30 3416.23 4715.16 7454.37 2.11e4
SparseLLM 10.96 16.57 13.38 19.07 28.77 19.29 2052.27 1536.51 538.61 289.17 687.48 677.13
OPT-6.7b
Sparsity Method BoolQ RTE HellaSwag WinoGrande ARC-e ARC-c OBQA Mean
Dense 66.12 56.03 50.49 65.27 65.72 30.63 27.60 51.69
70% SparseGPT 61.74 54.87 48.46 63.85 64.31 29.27 25.40 49.70
SparseLLM 60.61 54.51 48.8 62.9 64.14 30.03 26.60 49.66
80% SparseGPT 55.08 48.38 42.22 59.43 57.79 25.85 21.40 44.31
SparseLLM 58.69 51.26 43.78 59.67 58.38 26.88 22.00 45.81
90% SparseGPT 38.53 53.07 26.00 48.07 26.81 21.67 14.40 32.65
SparseLLM 46.48 52.71 26.21 51.70 27.44 19.71 13.40 33.95
3:4 SparseGPT 46.70 54.15 28.82 51.07 32.45 18.17 15.40 35.25
SparseLLM 53.49 53.42 36.24 53.51 43.94 22.61 17.40 40.09
OPT-13b
Sparsity Method BoolQ RTE HellaSwag WinoGrande ARC-e ARC-c OBQA Mean
Dense 65.87 57.76 52.44 66.02 67.82 33.46 28.62 53.14
70% SparseGPT 63.03 54.87 50.89 65.43 67.47 32.85 26.40 51.56
SparseLLM 63.85 55.23 50.73 65.67 66.46 31.83 27.20 51.57
80% SparseGPT 59.72 52.35 46.82 61.48 62.50 31.23 21.80 47.99
SparseLLM 60.89 53.07 46.19 62.12 62.21 30.38 23.00 48.27
90% SparseGPT 47.49 52.71 33.17 51.54 39.98 21.33 17.80 37.72
SparseLLM 53.43 52.71 38.19 52.96 46.68 25.26 17.40 40.95
3:4 SparseGPT 47.55 53.43 31.30 50.20 37.63 22.53 17.60 37.18
SparseLLM 51.13 52.35 38.51 55.96 49.24 24.83 21.40 41.92